OpenStudy (anonymous):

$29152000=$600(1+r)^92 solve:

ganeshie8 (ganeshie8):

use ur calculator

OpenStudy (anonymous):

Can you tell me first step?

ganeshie8 (ganeshie8):

divide 600 both sides

OpenStudy (anonymous):

I did that

OpenStudy (anonymous):

\((1+r)^{92}=(1+(92r)/92)^{92}\approx e^{92r}\) so the answer is roughly \(r\approx\frac1{92}\log(145760/3)\)